OpenStudy (anonymous):

LAST QUESTION AAHH :D Determine the type of boundary line and shading for the graph of the inequality 5x + y greater than or equal to -10. Dashed line with shading on the side that includes the origin Solid line with shading on the side that does not include the origin Dashed line with shading on the side that does not include the origin Solid line with shading on the side that includes the origin

OpenStudy (anonymous):

Medal and Fan awarded :)

OpenStudy (anonymous):

@iGreen i hope you could help :))

OpenStudy (igreen):

\(\ge\) or \(\le\) = solid line \(<\) or \(>\) = dashed line

OpenStudy (igreen):

\(>\) or \(\ge\) = shading that does not include the origin \(<\) or \(\le\) = shading that does include the origin

OpenStudy (igreen):

So what do you think? @CoinBank77